# Quasicrystalline 30 twisted bilayer graphene: Fractal patterns and electronic localization properties

### v1

@@ III.3 Suppression of the DOS
To understand how the suppression of the electronic density of states (DOS) is distributed in real space, we can utilize the geometric model introduced in Section II.1. This model allows us to identify the atomic sites with the strongest interlayer coupling between the two graphene layers.
Figures 7a to 7f show the emergent suppressions in the LDOS of the $30^{\circ}$-TBG systems. These suppressions are quantified by $S_{p}\equiv 1-\rho(\epsilon_{i},d)/\rho_{\rm max}(\epsilon_{i},d)$ at the energy $\epsilon_{i}=-0.536$ eV, that lies between the $\gamma$ and $\mu$-peak energies. A larger $S_{p}$ value indicates a greater suppression of the LDOS, suggesting stronger interlayer coupling between the graphene sheets. The minimum and maximum LDOS values correspond to $\rho_{\rm min}(\epsilon_{i},r_{2})=0.019$ a.u. and $\rho_{\rm max}(\epsilon_{i},c_{1})=0.05$ a.u., respectively.
@@
In Figure 7a, the black circle corresponds to a zoomed-in view of the structure close to the twist axis. The real-space distribution of suppression strengths exhibits a perfect 12-fold symmetry. The red circle, Fig. 7b, which corresponds to the dodecagonal centers, shows a LDOS distribution that only slightly deviates from 12-fold symmetry. In contrast, the green circle, corresponding to the vertices of different Stampfli tiles, exhibits a much stronger imbalance in the second ring closest to the center, violating the 12-fold symmetry. Figs. 7d to 7f illustrate the distributions of suppression strengths within the internal structures of Stampfli tiles, exhibiting double, triple and quadruple symmetry. Fig. 7g reveals that the emergent suppression strengths are highly localized at radial distances $r_{2}$ and $c_{2}$.
@@ III.4 Landau levels spectrum
Figure 8a shows a series of DOS for a pristine $30^{\circ}$-TBG system subjected to different perpendicular magnetic fields $B$. The LLs are broadened using $\eta=5$ meV [35]. The curves are offset in the $y$-axis for clarity. The spectra consist of a sequence of well-defined LL peaks labeled by the index $N$ with increasing field. Both the LL peak spacing and LL peak heights (related to the LL degeneracy) grow with increasing magnetic fields [46].
@@
The LL energies, $\epsilon_{N}$, are determined by using a smaller regularization parameter, $\eta=0.1$ meV, and fitting the computed DOS of each LL peak with a Lorentzian distribution. Figure 8b show that our results agree with the relation $\epsilon_{N}-\epsilon_{F}={\rm Sgn}(N)v_{F}\sqrt{2e\hslash|N|B}$, where $v_{F}$ the Fermi velocity, that is unique to massless Dirac fermions. For clarity, we subtract the Fermi energy $\epsilon_{F}$, measured at the energy of the charge neutrality point. Fitting both electron and hole branches of the experimental data yields a Fermi velocity of $v_{F}=(1.100\pm 0.001)\times 10^{6}$ m/s, as reported in Ref. [16]. In summary, our numerical results show a good agreement with the STS measurements, confirming massless Dirac fermions in $30^{\circ}$-TBG systems.
Let us compare the $30^{\circ}$-TBG LL energies with those of graphene monolayers. Figure 8c shows $\epsilon_{N}$ for $30^{\circ}$-TBG and graphene monolayer systems as a function of $(|N|B)^{1/2}$ for $25$ T. We verify that for low energies, corresponding to $|N|\lesssim 20$, the LL energies show, to very good accuracy, a linear dependence with $(|N|B)^{1/2}$. Thus, we conclude that the electronic properties of the decagonal quasicrystal in $30^{\circ}$-TBG behave like a system with two separated monolayer graphene in the low-energy region.
Finally, let us examine the large-$|N|$ limit, where one expects the LLs energies to deviate from the linear behavior [46, 35]. In the case of $30^{\circ}$-TBGs, these deviations arise due to long-range hopping terms that fit the nonlinear features of the dispersion relation [35, 47]. The inset of the Figure 8c gives the LL energy difference between $30^{\circ}$-TBG and graphene monolayer systems, $\Delta\epsilon_{N}=\epsilon^{\rm 30^{\circ}TBG}_{N}-\epsilon^{\rm G}_{N}$. We observe a notable discrepancy in the valence band of $\sim 20$ meV for $N=-15$. Nonetheless, the discrepancy is relatively small, reflecting the similarity of the DOS in the $30^{\circ}$-TBG and graphene systems at $B=0$. At higher energies, near $\epsilon\approx-1$ eV where the electronic states are localized (at $B=0$), we anticipate larger deviations. However, given that the LLs spacing at such energies is $\sim 10$ meV at $B=25$ T, experimental detection of these effects is challenging due to the broadening caused by intrinsic disorder and the significant excitation and/or doping required.
@@ IV Conclusions and discussion
We have studied the local spectral properties and emergent real-space localization in dodecagonal quasicrystal structures of $30^{\circ}$-TBG systems. The HHK method’s efficiency in calculating local spectral functions without relying on periodic boundary conditions proved invaluable for this study.
@@
We developed a simple emergent localization model for $30^{\circ}$-TBG systems based on two lattice construction methods. The emergent states associated with the singularities in the valence band are all identified with the Stampfli tiles, primarily those with imbalances in the non-perfect rings centered on the polygon vertices. Perfect and near-perfect rings are identified with decagonal centers that exhibit emergent localization measures with pronounced 12-fold symmetry and optimal balance.
We find that the most pronounced effects of interlayer coupling between two graphene layers of the $30^{\circ}$-TBG system are localized in the second ring closest to the center of the Stampfli square. We find that the most pronounced effects of interlayer coupling between two graphene layers of the $30^{\circ}$-TBG system are localized in the second ring centered on the polygon vertices ($r_{2}$) and in the centers of the Stampfly squares ($c_{2}$). This observation suggests a significant suppression of the LDOS at the intersection between the original and mirrored Dirac cones.

### v2

@@ III.3 Suppression of the DOS
We will now focus on the electronic density of states (DOS) suppression reported in Refs. [11, 18]. Our goal is to use the geometric model introduced in Section II.1 to understand the spatial distribution of the electronic states corresponding to the suppressed DOS.
Figures 7(a)-(f) show the emergent suppressions in the LDOS of the $30^{\circ}$-TBG systems. These suppressions are quantified by $S_{p}\equiv 1-\rho(\epsilon_{i},d)/\rho_{\rm max}(\epsilon_{i},d)$ at the energy $\epsilon_{i}=-0.536$ eV, that lies between the $\gamma$ and $\mu$-peak energies. A larger $S_{p}$ value indicates a greater suppression of the LDOS, suggesting stronger interlayer coupling between the graphene sheets. The minimum and maximum LDOS values correspond to $\rho_{\rm min}(\epsilon_{i},r_{2})=0.019$ a.u. and $\rho_{\rm max}(\epsilon_{i},c_{1})=0.05$ a.u., respectively.
@@
In Figure 7(a), the black circle corresponds to a zoomed-in view of the structure close to the twist axis. The real-space distribution of suppression strengths exhibits a perfect 12-fold symmetry. The red circle, Fig. 7(b), which corresponds to the dodecagonal centers, shows a LDOS distribution that only slightly deviates from 12-fold symmetry. In contrast, the green circle, corresponding to the vertices of different Stampfli tiles, exhibits a much stronger imbalance in the second ring closest to the center, violating the 12-fold symmetry. Figs. 7(d)-(f) illustrate the distributions of suppression strengths within the internal structures of Stampfli tiles, exhibiting double, triple and quadruple symmetry. Fig. 7(g) reveals that the emergent suppression strengths are highly localized at radial distances $r_{2}$ and $c_{2}$.
@@ III.4 Landau levels spectrum
Figure 8(a) shows a series of DOS for a pristine $30^{\circ}$-TBG system subjected to different perpendicular magnetic fields $B$. The LLs are broadened using $\eta=5$ meV [37]. The curves are offset in the $y$-axis for clarity. The spectra consist of a sequence of well-defined LL peaks labeled by the index $N$ with increasing field. Both the LL peak spacing and LL peak heights (related to the LL degeneracy) grow with increasing magnetic fields [48].
@@
The LL energies, $\epsilon_{N}$, are determined by using a smaller regularization parameter, $\eta=0.1$ meV, and fitting the computed DOS of each LL peak with a Lorentzian distribution. Figure 8(b) shows that our results agree with the relation $\epsilon_{N}-\epsilon_{F}={\rm Sgn}(N)v_{F}\sqrt{2e\hslash|N|B}$, where $v_{F}$ the Fermi velocity, that is unique to massless Dirac fermions. For clarity, we subtract the Fermi energy $\epsilon_{F}$, measured at the energy of the charge neutrality point. Fitting both electron and hole branches of the experimental data yields a Fermi velocity of $v_{F}=(1.100\pm 0.001)\times 10^{6}$ m/s, as reported in Ref. [18]. In summary, our numerical results show a good agreement with the STS measurements, confirming massless Dirac fermions in $30^{\circ}$-TBG systems.
Let us compare the $30^{\circ}$-TBG LL energies with those of graphene monolayers. Figure 8(c) shows $\epsilon_{N}$ for $30^{\circ}$-TBG and graphene monolayer systems as a function of $(|N|B)^{1/2}$ for $25$ T. We verify that for low energies, corresponding to $|N|\lesssim 20$, the LL energies show, to very good accuracy, a linear dependence with $(|N|B)^{1/2}$. Thus, we conclude that the electronic properties of the decagonal quasicrystal in $30^{\circ}$-TBG behave like a system with two separated monolayer graphene in the low-energy region.
Finally, let us examine the large-$|N|$ limit, where one expects the LLs energies to deviate from the linear behavior [48, 37]. In the case of $30^{\circ}$-TBGs, these deviations arise due to long-range hopping terms that fit the nonlinear features of the dispersion relation [37, 49]. The inset of the Figure 8(c) gives the LL energy difference between $30^{\circ}$-TBG and graphene monolayer systems, $\Delta\epsilon_{N}=\epsilon^{\rm 30^{\circ}TBG}_{N}-\epsilon^{\rm G}_{N}$. We observe a notable discrepancy in the valence band of $\sim 20$ meV for $N=-15$. Nonetheless, the discrepancy is relatively small, reflecting the similarity of the DOS in the $30^{\circ}$-TBG and graphene systems at $B=0$. At higher energies, near $\epsilon\approx-1$ eV where the electronic states are localized (at $B=0$), we anticipate larger deviations. However, given that the LLs spacing at such energies is $\sim 10$ meV at $B=25$ T, experimental detection of these effects is challenging due to the broadening caused by the intrinsic disorder and the significant excitation and/or doping required.
@@ IV Conclusions and discussion
We have studied the local spectral properties and the emergence of real-space electronic localization in dodecagonal quasicrystal structures of $30^{\circ}$-TBG systems. The HHK method’s efficiency in calculating local spectral functions without relying on periodic boundary conditions outperforms atomistic calculations [15] by orders of magnitude. Furthermore, our findings exhibit excellent agreement with the localized states reported in Ref. [15] and the experimental density of states observed in Refs. [11, 18].
@@
We developed a simple geometric model, applicable to the two lattice construction methods, to analyze the LDOS features in $30^{\circ}$-TBG systems. The emergent states associated with the singularities in the valence band are all identified with the Stampfli tiles, primarily those with imbalances in the non-perfect rings centered on the polygon vertices. Perfect and near-perfect rings are identified with decagonal centers that exhibit emergent localization measures with pronounced 12-fold symmetry and optimal balance.
We find that the most pronounced effects of interlayer coupling between two graphene layers of the $30^{\circ}$-TBG system are localized in the second ring centered on the polygon vertices ($r_{2}$) and in the centers of the Stampfly squares ($c_{2}$). This observation suggests a significant suppression of the LDOS at the intersection between the original and mirrored Dirac cones.
As an outlook, the efficiency of our method enables the analysis of twisted multilayer graphene systems [9, 10], both commensurate and incommensurate, which involve a significantly larger number of atomic sites. Of particular interest are the recently observed moiré quasicrystals in twisted trilayer graphene [50], which present a promising direction for further development of our work.